# Smaller and More Flexible Cuckoo Filters

### v1

@@ 4.1 Just-in-time compilation
We have implemented the improved Cuckoo filters for integer keys in just-in-time compiled Python using the numba package [18] with typed numpy arrays. This offers several benefits, such as highly optimized machine code, the use of LLVM intrinsics, and the option to choose parameters during runtime, but before compilation. For example, this allows us to provide all parameters of the hash functions as compile-time constants for additional optimizations, achieving speeds comparable to compiled C or C++ code, sometimes even faster, due to the increased possibilities for optimization. Code is available at https://gitlab.com/rahmannlab/cuckoo-filters.
@@ 4.2 Distinguishing full slots from empty slots
We have so far not discussed how we can distinguish an empty slot from a full slot. The hash function $f_{0}:\mathcal{U}\to[2^{k}]$ maps a key $x$ to its $k$-bit fingerprint $f_{0}(x)$. We use one of the possible fingerprint values (00) to indicate an empty slot. Hence, we reduce the set of valid fingerprints and instead use a hash function $f^{\prime}_{0}:\mathcal{U}\to[2^{k}-1]$. The fingerprint is given by $f_{0}(x):=f^{\prime}_{0}(x)+1$. The loss of one fingerprint value leads to a higher FPR of $1/(2^{k}-1)$, the probability that two random fingerprints collide. Although this effect is measurable for small $k$, it can be considered as negligible for typically used larger values of $k$; see also Figure 4 in Section 5.2. The same compromise was made for the original Cuckoo filter [11].
@@ 4.3 Bit-parallel engineering
We load the contents of a window into a 64-bit integer register that we (conceptually) partition into four adjacent slots (and some unused bits). To check whether there is an empty slot (all zeros), we use two pre-computed masks, lo and hi = lo << (q-1), indicating the least significant and most significant bits in each slot, respectively. By computing the bit pattern e := (x - lo) & (~x) & hi, the 1-bits of e are those 1-bits of hi that indicate which slots are empty. If e is non-zero, there exists an empty slot and its index (from the right) can be computed by dividing the number of trailing zeros of e (obtained by a cttz CPU instruction) by q. In the following example, the result is nonzero and has 7 trailing zeros, indicating that slot 0 in $x$ is empty. Comparing all 4 slots at the same time in this bit-parallel manner saves the overhead of a loop or separate comparisons.
@@
            x - lo = 10001101 | $11$ | 11100100 | $11$   -: subtraction
                ~x = 01110000 | $11$ | 00011001 | $11$   ~: bit-wise inversion
       (x-lo) & ~x = $00$ | $11$ | $00$ | $11$   &: bit-wise and
  (x-lo) & ~x & hi = $00$ | 10000000 | $00$ | 10000000   result e, 7 trailing 0s
Similarly, to search for a fingerprint in any of the four slots, we can create a single bit mask that contains the valid bit patterns for each slot. For $(2,4)$ windowed Cuckoo filters, we use one choice bit and two window offset bits. Assume that the 5-bit fingerprint is 11100, that we are examining window choice 1, and that the offset bits increase with decreasing slot number. Then the valid slot bit patterns can be encoded in a mask y, and we can check if the bit patterns agree in one of the slots (and which one) by taking the bit-wise XOR x^y and checking for a zero slot like above. In the following example, we find the fingerprint in slot 1.
@@
       x = 10001111 | $00$ | 11011100 | $00$   window contents x
   x ^ y = 00010011 | 10111100 | $00$ | 11111100   ^: XOR
@@ 5 Evaluation
We start by comparing bucketed with windowed Cuckoo filters. We first evaluate the time-memory trade-off for increasing loads (Section 5.1). In Section 5.2, we evaluate the actual FPRs, loads and overhead factors for different desired FPRs at target loads of $0.98\cdot T$, where $T$ is the theoretical load threshold for a given configuration (Table 2). More technical aspects, such as achievable loads, depending on the maximum random walk length during insertion or on the number of inserted keys, are given in Appendix C. Performance benchmarks of our implementation are shown in Section 5.3.
@@
We evaluate the original Vector Quotient filter implementation from [28] (downloaded from https://github.com/splatlab/vqf, written in C++) and the original Prefix filter implementation from [10] (from https://github.com/TomerEven/Prefix-Filter, written in C++) using a Cuckoo filter or VQF as a spare.
@@
Evaluations were run on a PC workstation with an AMD Ryzen 9 9950X CPU with 16 cores and hyperthreading and 64 GB of DDR5 memory (6000 MHz, CL40). All benchmarks were performed on random unsigned 64-bit integer keys, generated with numpy. Reported times are wall times, including just-in-time compilation and excluding data load time, except for multi-threaded insertion, where reader and inserter threads load and insert keys in parallel (see Appendix A). We report averages over five runs.
@@ 5.2 Actual overhead factors, loads and FPRs
Figure 4: Comparison of actual properties (memory overhead, actual load, empirical FPR) of Cuckoo filter types (distinguished by color) designed for FPRs of $2^{-k}$ for varying $k$ (x-axis), using a fixed maximum random walk length of 10 000, with $n=2\times 10^{9}$ keys and 5 subfilters for parallelization. Insertions are performed only if the filter does not already report a key as present. Left: Actual memory overhead factor $C$ for a Cuckoo filter with $Cnk$ bits for a target FPR of $2^{-k}$; lower is better. Middle: Empirical loads. Right: Empirical FPRs, relative to the target FPR of $2^{-k}$, as log ratios. Points below the horizontal red line at 0.00 have an actual FPR that is better than the target FPR.
According to the observations of Appendix C, we choose a filter size that ensures a load of at most 98% of the theoretical load threshold and a maximum walk length of 10 000 steps. Using $n=2\cdot 10^{9}$ random 64-bit integer keys, we now compare the resulting actual memory overhead factors, empirical FPRs and observed load factors for different values of $k$ (Figure 4). The actual memory overhead factor is obtained by dividing the actual number of used bits by $nk$. The actual load is the number of occupied slots, divided by the number of total slots. The empirical FPR is obtained by querying $n^{\prime}=2\cdot 10^{9}$ random keys that were not previously inserted into the filter, computed as the number of times the lookup erroneously returned True, divided by $n^{\prime}$.
@@
For all Cuckoo filter variants, the overhead factor decreases for increasing $k$ (i.e., for lower FPRs; see Table 1 and Figure 4 left). Windowed Cuckoo filters with window size 2 have the lowest overhead factors, since a window size of 2 saves one bit per slot compared to buckets or windows of size 4, and they achieve high load thresholds. Although buckets of size 2 use the same number of bits per slot, their load threshold is much lower, and they therefore have higher overhead factors compared to windows of size 2. For small $k$, buckets of size 2 have a lower overhead factor compared to windows of size 4, but from $k\geq 8$, the overhead factor for windows of size 4 is lower, since the higher load compensates for the additional bit per slot. Cuckoo filters with buckets of size $4$ have higher overhead factors compared to windows of size $4$ due to the load threshold.
@@
We target an FPR of $2^{-k}$ by using $k+2$ or $k+3$ bits per slot, for a window or bucket size of 2 or 4, respectively. In practice, several other effects increase or decrease the actual FPR. Figure 4 (right) shows log-ratios between empirical FPR and $2^{k}$. For $k\geq 5$, the observed FPR is lower than $2^{-k}$, because the filter is not full. Actual non-present keys have a small chance of hitting an empty slot instead of a stored fingerprint, decreasing the FPR. Since the load decreases for small $k$ (see above), the observed FPR should also decrease for small $k$. However, for $k\leq 4$, the window-based filters show a comparably high FPR. Here we see the effect of sacrificing one fingerprint value to indicate empty slots, e.g. for $k=3$, we do not have 8 but only 7 usable fingerprints. Since for the windowed versions, the fingerprint size is only $k$, this effect is severe. For the bucketed versions, we have 1 or 2 additional fingerprint bits (for bucket sizes 2 or 4, respectively) and no window offset bits, so the effect is less severe. While the interplay of these different effects on the FPR for small $k$ is interesting, it is mostly irrelevant in practice, as Cuckoo filters are most useful for $k\geq 9$.
@@ 5.3 Performance Benchmarks
The speedup factor from parallelization is almost linear, both for insertions and for queries for up to 5 threads (Figure 5 bottom, see Appendix A for details). For more threads, the distributing main thread starts to become the bottleneck. The speedup for insertions is slightly lower than for queries, probably due to higher memory bandwidth utilization.
@@ 5.4 Comparison with Prefix Filters and Vector Quotient Filters
    Both the VQF and the Prefix filter only support a FPR of $\approx 2^{-8}$ (and $2^{-16}$ for the VQF) and their performance is highly optimized for this special case. A comparable Cuckoo filter with 8 bits per slot only has an FPR of $2^{-5}$. In contrast, we support all FPRs with different optimizations due to the advantages of just-in-time compilation, but at the cost of being less optimized for one special case.
@@
    The Prefix filter implementation only works on CPUs that support AVX512 instructions; there is no fallback implementation.
@@ 6 Conclusion and Discussion
We introduced improvements of Cuckoo filters, enabling their more flexible use and reducing their overhead factor from $1.05(1+3/k)$ to $1.06(1+2/k)$, switching from a $(2,4)$ bucketed layout to a $(2,2)$ windowed layout. This is a practical improvement for relevant FPRs of $2^{-k}$ with $9\leq k\leq 30$ (and in theory up to $k\leq 102$). In comparison to other state-of-the-art filters, $(2,2)$ windowed Cuckoo filters have the smallest space overhead and competitively fast lookup times, especially for optimized cases, such as $k=14$, where four 16-bit slots can be examined in a bit-parallel manner in a single 64-bit integer. In that case, the overhead factor of $1.20$ even outperforms static XOR filters [15] that need a factor of $1.23$, originally advertised as “faster and smaller than Bloom and Cuckoo filters”. (Of course, by now, better static filters exist [8].) Low space requirements and fast query times at low FPRs make the improved Cuckoo filters valuable candidates for many applications, e.g., in DNA sequence analysis applications [32], where large exact hash tables may be replaced by smaller filters at the cost of a few false positive queries.
For larger FPRs with $k\leq 8$, better alternatives exist. Especially for $k=8$, Prefix filters are highly optimized concerning fast insert, where Cuckoo filters are not competitive. For even lower values of $k$, Blocked Bloom filters are simpler and can be competitive.
Another interesting aspect of this work is that just-in-time compiled Python code can adjust to run-time specified parameters (such as $k$). The compiler can use them as compile-time constants, thereby optimizing the code differently for different parameters, outperforming even hand-crafted code. The obtained lookup performance of 51.7 million keys per second in $(2,2)$ bucketed Cuckoo filters at an FPR of $2^{-14}$ and an overhead factor of 1.28 outperforms the hand-crafted AVX512 code of Prefix filters at an FPR of $2^{-8}$ that achieves “only” 44.6 million keys per second at an overhead factor of 1.45.
@@
Appendix A Parallelization
We parallelize the filter as follows: We create $F$ subfilters, such that the total number of buckets (or windows) $B$ is divided equally among the $F$ subfilters, with $\lceil B/F\rceil$ buckets or windows per subfilter. Each subfilter is a cache-line aligned bit array. For inserting or querying a key, the subfilter index is computed by an additional hash function $f_{\text{sub}}:\mathcal{U}\mapsto[F]$. The bucket (or window) address hash function now maps the key $x$ to any of the $\lceil B/F\rceil$ buckets (or windows) in the subfilter. To insert keys into the subfilter, one main thread delegates the work to one thread per subfilter, responsible for performing the subfilter insertion. To avoid locks, the threads communicate via message buffers, as in [33]. For querying keys, all threads only need read access and hence no synchronization between threads is needed. The queries are thus distributed evenly across all threads.
Appendix B Hash functions
@@
Appendix C Comparison of load thresholds with achievable loads
We created datasets of random 64-bit integers and counted the number of inserted keys until the insertion fails for the first time. We tabulated the percentage of filled and empty slots in the filter to compute the achievable load of bucketed and windowed Cuckoo filters with different parameters. The reported results are averaged over five runs.

### v2

@@ 4.1 Just-in-time compilation
We have implemented the improved Cuckoo filters for integer keys in just-in-time compiled Python using the numba package [19] with typed numpy arrays. This offers several benefits, such as highly optimized machine code, the use of LLVM intrinsics, and the option to choose parameters during runtime, but before compilation. For example, this allows us to provide all parameters of the hash functions as compile-time constants for additional optimizations, achieving speeds comparable to compiled C or C++ code, or sometimes even faster, due to the increased possibilities for optimization. Since most state-of-the-art filters are implemented in compiled languages, we compared our reference implementation of the original Cuckoo filter with an existing C++ implementation [12]; see Appendix D. Since both implementations achieve similar running times, using just-in-time compiled Python comes with little penalties and allows us to compare our implementation with other state-of-the-art filters written in low-level languages. The code is available at https://gitlab.com/rahmannlab/cuckoo-filters.
@@ 4.2 Distinguishing full slots from empty slots
We have so far not discussed how we can distinguish an empty slot from a full slot. The hash function $f_{0}:\mathcal{U}\to[2^{k}]$ maps a key $x$ to its $k$-bit fingerprint $f_{0}(x)$. We use one of the possible fingerprint values (by convention 0) to indicate an empty slot. Hence, we reduce the set of valid fingerprints and instead use a hash function $f^{\prime}_{0}:\mathcal{U}\to[2^{k}-1]$. The fingerprint is given by $f_{0}(x):=f^{\prime}_{0}(x)+1$. The loss of one fingerprint value leads to a higher FPR of $1/(2^{k}-1)$, the probability that two random fingerprints collide. Although this effect is measurable for small $k$, it can be considered as negligible for typically used larger values of $k$; see also Figure 4 in Section 5.2. The same compromise was made for the original Cuckoo filter [12].
@@ 4.3 Bit-parallel engineering
We load the contents of a window into a 64-bit integer register that we (conceptually) partition into four adjacent slots (and some unused bits). To check whether there is an empty slot (all zeros), we use two pre-computed masks, lo and hi = lo << (q-1), indicating the least significant and most significant bits in each slot, respectively. By computing the bit pattern e := (x - lo) & (˜x) & hi, the 1-bits of e are those 1-bits of hi that indicate which slots are empty. If e is non-zero, there exists an empty slot and its index (from the right) can be computed by dividing the number of trailing zeros of e (obtained by a cttz CPU instruction) by q. In the example , the result is nonzero and has 7 trailing zeros, indicating that slot 0 in $x$ is empty. Comparing all 4 slots at the same time in this bit-parallel manner saves the overhead of a loop or separate comparisons.
@@
            x - lo = 10001101 | 11111111 | 11100100 | 11111111   -: subtraction
                ~x = 01110000 | 11111111 | 00011001 | 11111111   ~: bit-wise inversion
       (x-lo) & ~x = 00000000 | 11111111 | 00000000 | 11111111   &: bit-wise and
  (x-lo) & ~x & hi = 00000000 | 10000000 | 00000000 | 10000000   result e, 7 trailing 0s
Similarly, to search for a fingerprint in any of the four slots, we can create a single bit mask that contains the valid bit patterns for each slot. For $(2,4)$ windowed Cuckoo filters, we use one choice bit and two window offset bits. Assume that the 5-bit fingerprint is 11100, that we are examining window choice 1, and that the offset bits increase with decreasing slot number. Then the valid slot bit patterns can be encoded in a mask y, and we can check if the bit patterns agree in one of the slots (and which one) by taking the bit-wise XOR xˆy and checking for a zero slot like above. In the following example, we find the fingerprint in slot 1.
@@
       x = 10001111 | 00000000 | 11011100 | 00000000   window contents x
   x ^ y = 00010011 | 10111100 | 00000000 | 11111100   ^: XOR
4.4 Parallelization
We parallelize the filter as follows: We create $F$ independent subfilters, such that the total number of buckets (or windows) $B$ is divided equally among the $F$ subfilters, with $\lceil B/F\rceil$ buckets or windows per subfilter. Each subfilter is a cache-line-aligned bit array. For inserting or querying a key, the subfilter index is computed by an additional hash function $f_{\text{sub}}:\mathcal{U}\mapsto[F]$. The bucket (or window) address hash function now maps the key $x$ to any of the $\lceil B/F\rceil$ buckets (or windows) in the subfilter. To insert keys into the subfilter, one main thread delegates the work to one thread per subfilter, responsible for performing the subfilter insertion. To avoid locks, the threads communicate via message buffers using a consumer-producer principle, as in [34]. The main thread computes the subfilter number $f_{\text{sub}}(x)$ for each key $x$ and copies the key to a buffer assigned to this subfilter. For each subfilter, several such buffers exist, and these buffers are only read by one consumer thread responsible for insertions into this particular subfilter. If a buffer is full, the main thread sets an atomic volatile ready_to_read flag and continues inserting keys into other buffers. Among the buffers assigned to a specific subfilter, the corresponding consumer thread searches for a buffer marked as ready_to_read and inserts the keys into the subfilter. After all keys in the buffer are inserted, the consumer thread marks the buffer as ready_to_write indicating to the main thread that it can again insert keys into the buffer.
Assuming that queries are processed once all keys are inserted, and given that queries only need read access, it follows that no synchronization between query threads is needed. Queries are thus distributed evenly across all threads.
@@ 5 Evaluation
We start by comparing bucketed with windowed Cuckoo filters. We first evaluate the time-memory trade-off for increasing loads (Section 5.1). In Section 5.2, we evaluate the actual FPRs, loads and overhead factors for different desired FPRs at target loads of $0.98\cdot T$, where $T$ is the theoretical load threshold for a given configuration (Table 2). More technical aspects, such as achievable loads, depending on the maximum random walk length during insertion or on the number of inserted keys, are given in Appendix B. Performance benchmarks of our implementation are shown in Section 5.3.
@@
We evaluate the original Vector Quotient filter implementation from [29] (downloaded from https://github.com/splatlab/vqf, written in C++) and the original Prefix filter implementation [11] (from https://github.com/TomerEven/Prefix-Filter, written in C++) using a Cuckoo filter or VQF as a spare.
@@
Evaluations were run on a PC workstation with an AMD Ryzen 9 9950X CPU with 16 cores and hyperthreading and 64 GB of DDR5 memory (6000 MHz, CL40). All benchmarks were performed on random unsigned 64-bit integer keys, generated with numpy. Reported times are wall times, including just-in-time compilation and excluding data load time, except for multi-threaded insertion, where reader and inserter threads load and insert keys in parallel. We report averages over five runs.
@@ 5.2 Actual overhead factors, loads and FPRs
Figure 4: Comparison of actual properties (memory overhead, actual load, empirical FPR) of Cuckoo filter types (distinguished by color) designed for FPRs of $2^{-k}$ for varying $k$ (x-axis), using a fixed maximum random walk length of 10 000, with $n=2\times 10^{9}$ keys and 5 subfilters for parallelization. Insertions are performed only if the filter does not already report a key as present. Left: Actual memory overhead factor $C$ for a Cuckoo filter with CnkCnk bits for a target FPR of $2^{-k}$; lower is better. Middle: Empirical loads. Right: Empirical FPRs, relative to the target FPR of $2^{-k}$, as log ratios. Points below the horizontal red line at 0.00 have an actual FPR that is better than the target FPR.
According to the observations of Appendix B, we choose a filter size that ensures a load of at most 98% of the theoretical load threshold and a maximum walk length of 10 000 steps. However, the actual maximum random walk length is usually much lower; see Appendix C. Using $n=2\cdot 10^{9}$ random 64-bit integer keys, we now compare the resulting actual memory overhead factors, empirical FPRs and observed load factors for different values of $k$ (Figure 4). The actual memory overhead factor is obtained by dividing the actual number of used bits by $nk$. The actual load is the number of occupied slots, divided by the number of total slots. The empirical FPR is obtained by querying $n^{\prime}=2\cdot 10^{9}$ random keys that were not previously inserted into the filter, computed as the number of times the lookup erroneously returned True, divided by $n^{\prime}$.
@@
For all Cuckoo filter variants, the overhead factor decreases for increasing $k$ (i.e., for lower FPRs; see Table 1 and Figure 4 left). Windowed Cuckoo filters with window size 2 have the lowest overhead factors, since a window size of 2 saves one bit per slot compared to buckets or windows of size 4, and they achieve high load thresholds. Although buckets of size 2 use the same number of bits per slot, their load threshold is much lower, and they therefore have higher overhead factors compared to windows of size 2. For small $k$, buckets of size 2 have a lower overhead factor compared to windows of size 4, but from $k\geq 8$, the overhead factor for windows of size 4 is lower, since the higher load compensates for the additional bit per slot. Cuckoo filters with buckets of size $4$ have higher overhead factors compared to windows of size $4$ due to the higher load threshold.
@@
We target an FPR of $2^{-k}$ by using $k+2$ or $k+3$ bits per slot, for a window or bucket size of 2 or 4, respectively. In practice, several other effects increase or decrease the actual FPR. Figure 4 (right) shows log-ratios between empirical FPR and $2^{-k}$. For $k\geq 5$, the observed FPR is lower than $2^{-k}$, because the filter is not full. Actual non-present keys have a small chance of hitting an empty slot instead of a stored fingerprint, decreasing the FPR. Since the load decreases for small $k$ (see above), the observed FPR should also decrease for small $k$. However, for $k\leq 4$, the window-based filters show a comparably high FPR. Here we see the effect of sacrificing one fingerprint value to indicate empty slots, e.g. for $k=3$, we do not have 8 but only 7 usable fingerprints. Since for the windowed versions, the fingerprint size is only $k$, this effect is severe. For the bucketed versions, we have 1 or 2 additional fingerprint bits (for bucket sizes 2 or 4, respectively) and no window offset bits, so the effect is less severe. While the interplay of these different effects on the FPR for small $k$ is interesting, it is not so important in practice, as most applications aim at FPRs of $2^{-k}$ with $k\gg 5$.
@@ 5.3 Performance Benchmarks
The speedup factor from parallelization is almost linear, both for insertions and for queries for up to 5 threads (Figure 5 bottom, see Appendix 4.4 for details). For more threads, the distributing main thread starts to become the bottleneck. The speedup for insertions is slightly lower than for queries, probably due to higher memory bandwidth utilization.
@@ 5.4 Comparison with Prefix Filters and Vector Quotient Filters
    Both the VQF and the Prefix filter only support a FPR of $\approx 2^{-8}$ (and $2^{-16}$ for the VQF), and their performance is optimized for this special case. A comparable Cuckoo filter with 8 bits per slot has an FPR of $2^{-5}$. In contrast, we support all FPRs with different optimizations due to the advantages of just-in-time compilation, but at the cost of being less optimized for one special case.
@@
    The available Prefix filter implementation only works on CPUs that support AVX512 instructions; there is no fallback implementation.
@@
Summary.
Our design has a smaller space overhead compared to state-of-the are filters, while lookup times remain competitive. If fast insertion times are essential, and an FPR of $2^{-8}$ is tolerable, Prefix filters are preferable.
@@ 6 Conclusion and Discussion
We introduced improvements of Cuckoo filters, enabling their more flexible use and reducing their overhead factor from $1.05(1+3/k)$ to $1.06(1+2/k)$, switching from a $(2,4)$ bucketed layout to a $(2,2)$ windowed layout. This is a practical improvement for relevant FPRs. For an FPR of $2^{-8}$, $(2,2)$ windowed Cuckoo filters need $\approx 10.6$ GB to index 10 billion keys, compared to $\approx 11.6$ GB for $(2,4)$ bucketed Cuckoo filters. Another interesting aspect of this work is that just-in-time compiled Python code can adjust to run-time specified parameters (such as $k$), resulting in a flexible implementation of Cuckoo filters that is competitive in speed with a C++ implementation.
In comparison to other state-of-the-art filters, our implementation of $(2,2)$ windowed Cuckoo filters has the smallest space overhead and competitively fast lookup times, especially for optimized cases, such as $k=14$, where four 16-bit slots can be examined in a bit-parallel manner in a single 64-bit integer. In that case, the overhead factor of $1.20$ even outperforms static XOR filters [16] that need a factor of $1.23$, originally advertised as “faster and smaller than Bloom and Cuckoo filters”. (Of course, by now, better static filters exist [9].)
For $k=8$, Prefix filters are highly optimized concerning fast insertions, here Cuckoo filters are not competitive in terms of speed. In this case, the increased speed may outweigh the slightly larger space overhead of Prefix filters. For even lower values of $k$, Blocked Bloom filters are simpler and can be competitive.
In summary, the low space requirements and fast query times make the improved Cuckoo filters valuable candidates for many applications, e.g., in DNA sequence analysis applications [33], where large exact hash tables may be replaced by smaller filters at the cost of a few false positive queries.
@@
Appendix A Hash functions
@@
Appendix B Comparison of load thresholds with achievable loads
We created datasets of random 64-bit integers and counted the number of inserted keys until the insertion fails for the first time. We tabulated the percentage of filled and empty slots in the filter to compute the achievable load of bucketed and windowed Cuckoo filters with different parameters. We show the average (main line), the minimum and maximum (shaded region) over five runs.